OpenStudy (anonymous):

does anyone know how to convert 6/5 into a decimal

TheSmartOne (thesmartone):

divide it in your calculator and you get 1.2

OpenStudy (anonymous):

Type 6 divided by 5 on your calculator. You should get 1.2

OpenStudy (anonymous):

thanks

TheSmartOne (thesmartone):

Or if you know\[\frac{ 1 }{ 5 }+\frac{ 1 }{ 5 }+\frac{ 1 }{ 5 }+\frac{ 1 }{5 }+\frac{ 1 }{ ? }\] then you know\[\frac{ 1 }{ 5 }=.2\] and therefore \[.2+.2+.2+.2+.2=1\]

TheSmartOne (thesmartone):

Ugh typo

TheSmartOne (thesmartone):

\[\frac{ 1 }{ 5 }+\frac{ 1 }{ 5 }+\frac{ 1 }{ 5 }+\frac{ 1 }{5 }+\frac{ 1 }{5 }=1\] not

TheSmartOne (thesmartone):

\[\frac{ 1 }{ 5 }+\frac{ 1 }{ 5 }+\frac{ 1 }{ 5 }+\frac{ 1 }{5 }+\frac{ 1 }{ ? }\]

OpenStudy (anonymous):

Or just divide the numerator by the denominator ;)

OpenStudy (anonymous):

so 6/5 is equal to 1.2 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you so much alright bye